How do I draw the graph of y = (1/2)sin(-x) if there are no points?
Do I solve for it? If so, how? Please explain the steps to me, I really don't understand this! Thank you!

Answer by lwsshak3(11628)

How do I draw the graph of y = (1/2)sin(-x) if there are no points?
**
Equation for the sin function: y=Asin(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
For given equation: y=(1/2)sin(-x)
Amplitude=1/2
B=1
Period=2π/1=2π
C=0
Phase shift: none
..
Graphing for one period:
Start with coordinates of basic sin curve: y=sin x (amplitude=1, period=2π)
(0,0), (π/2,1), (π,0), (3π/2,-1), (2π,0)
..
change amplitude to 1/2
coordinates: (0,0), (π/2,1/2), (π,0), (3π/2,-1/2), (2π,0)
..
change x to -x (curve reflects about the x-axis)
coordinates: (0,0), (π/2,-1/2), (π,0), (3π/2,1/2), (2π,0)
These are the coordinates of the given sin function which you can use to graph the function.